### v1

# MORDA: A Synthetic Dataset to Facilitate Adaptation of Object Detectors to Unseen Real-target Domain While Preserving Performance on Real-source Domain

Abstract
Deep neural network (DNN) based perception models are indispensable in the development of autonomous vehicles (AVs). However, their reliance on large-scale, high-quality data is broadly recognized as a burdensome necessity due to the substantial cost of data acquisition and labeling. Further, the issue is not a one-time concern, as AVs might need a new dataset if they are to be deployed to another region (real-target domain) that the in-hand dataset within the real-source domain cannot incorporate. To mitigate this burden, we propose leveraging synthetic environments as an auxiliary domain where the characteristics of real domains are reproduced. This approach could enable indirect experience about the real-target domain in a time- and cost-effective manner. As a practical demonstration of our methodology, nuScenes and South Korea are employed to represent real-source and real-target domains, respectively. That means we construct digital twins for several regions of South Korea, and the data-acquisition framework of nuScenes is reproduced. Blending the aforementioned components within a simulator allows us to obtain a synthetic-fusion domain in which we forge our novel driving dataset, MORDA: Mixture Of Real-domain characteristics for synthetic-data-assisted Domain Adaptation. To verify the value of synthetic features that MORDA provides in learning about driving environments of South Korea, 2D/3D detectors are trained solely on a combination of nuScenes and MORDA. Afterward, their performance is evaluated on the unforeseen real-world dataset (AI-Hub111This research (paper) used datasets from High-precision data collection vehicle daytime city road data. All data information can be accessed through AI-Hub (http://www.aihub.or.kr ).) collected in South Korea. Our experiments present that MORDA can significantly improve mean Average Precision (mAP) on AI-Hub dataset while that on nuScenes is retained or slightly enhanced. Details on MORDA will be available at https://morda-e8d07e.gitlab.io.
I INTRODUCTION
Deep-learning-based techniques have obtained growing attention and have become a major trend in challenging perception tasks of autonomous vehicles (AVs). Such dominance occurs across types of sensor modality: camera [1, 2, 3], LiDAR [4, 5, 6], RADAR [7], and even combination of those [8, 9, 10]. Despite their thrives in AVs, one inherent challenge is the dependency on a large-scale driving dataset. In general, building a dataset for perception tasks involves (1) equipping a car with the desired sensor suite, (2) deploying the car to the regions of interest to collect sensor data, and lastly, (3) generating the ground truth (GT) labels for sensor data according to labeling rules. As widely known, this entire process is time-consuming and cost-heavy.
This challenge gets even more severe when deploying AVs that have been developed for one domain (real-source domain, ) to another (real-target domain, ). In this context, one preferable choice for safety-critical AVs would be rebuilding the training dataset with GT labels for the novel domain. However, this method is indeed pricey.
To alleviate the labeling cost, research on unsupervised domain adaptation (UDA) has been actively conducted. UDA methods assume that only the is labeled, and they exploit the knowledge from the annotated for better adaptation (i.e., performance) of DNNs to the unlabeled without additional labeling costs [11, 12]. Nevertheless, dispatching the sensor-equipped vehicle to for raw sensor data collection is inevitable. Furthermore, the engineering cost for synchronizing raw data from sensors operating at different frequencies is non-negligible.
[b] Sensor setup BBox annotation Camera* Resolution LiDAR Channel Sensor suite reference DOF† Road obstacle‡ Frequency(Hz) VIPER [13] 1 19201080 ✗ ✗ - 9 ✗ ~15 VKITTI2 [14] 5 1242375 ✗ ✗ KITTI [15] 9 ✗ - SYNTHIA-AL [16] 1 640480 ✗ ✗ - 7 ✗ 25 SHIFT [17] 5 1200800 1 128 - 9 ✗ 10 MORDA (Ours) 6 1600900 1 32 nuScenes [18] 9 ✓ 20
Only multi-view cameras are counted, and stereo cameras are not included.
Traffic cones and barriers.
Relevant information is not available.
Degree of freedom, 7 and 9 indicates the orientation is represented with (yaw) and (roll, pitch, and yaw), respectively.
Recently, game engines and simulators have been increasingly employed to reduce the burden given that a large-scale dataset with accurate GT labels can be constructed in a time- and cost-efficient manner [19, 13, 17, 16, 20, 14]. However, many synthetic driving datasets at present are not specifically designed to replicate particular real-world domains, including driving environments, sensor suites, object categories, and labeling policies. Consequently, inconsistencies between those characteristics of real-world and synthetic datasets might further widen the gap between them, potentially resulting in performance degradation of DNNs when trained on real and synthetic datasets combined.
Given the mentioned constraints in UDA and existing synthetic datasets, the research objective of this paper is to develop a synthetic dataset that enhances the adaptation of DNNs to while their performance on is not degraded. Note that this scenario is challenging as (1) is completely unforeseen, meaning sensor data of the target domain is not used for training DNNs, and (2) the synthetic dataset should improve the generalization ability of DNNs as they need to perform well on both and .
We believe an adequate synthetic dataset for our purpose needs to be equipped with the following attributes; (1) Virtual version of geographical features of driving environments in the target domain (), which enables cost-effective indirect exposure of to DNNs. (2) Consistency with in terms of sensor suite and labeling process to keep the characteristics of created synthetic dataset paired with . In this regard, we propose a methodology to generate a synthetic dataset in a synthetic-fusion domain () where the aforementioned aspects of and are well integrated. As a demonstration, we construct a novel synthetic dataset, MORDA which takes nuScenes for and South Korea for . Next, we evaluate the efficacy of our MORDA dataset in a challenging scenario, as illustrated in Fig. 1. Firstly, 2D/3D object detectors are trained solely on nuScenes () combined with MORDA (). Afterward, they are evaluated on the unforeseen AI-Hub dataset collected in South Korea (). Our extensive experiments present that MORDA allows notable detection-performance gain on while successfully maintaining the performance on simultaneously. In short, the main contributions are:
We present MORDA, a synthetic dataset featuring the virtual version of the real driving environments in South Korea and replicating key components of nuScenes in data acquisition and labeling.
Using MORDA, we show that utilizing could be a financially-viable method for object detectors to adapt without performance degradation in .
II RELATED WORKS
II-A Object Detection
The field of object detection has shown surprising progress with DNNs in recent years, which aims to the accurate prediction of bounding box (BBox) to localize and classify objects within the sensor’s field-of-view. The prosperity of DNN-based object detection applies to not only 2D but also challenging 3D detection tasks across various sensor modalities [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, the inherent challenge of learning algorithms remains that performance degradation occurs when DNNs are evaluated on a new domain which is unforeseen during training [21, 17]. Note that addressing such is imperative for AVs whose driving environments or conditions can often change, e.g., deploying the developed AVs to another country ().
II-B Unsupervised Domain Adaptation (UDA)
To tackle the performance drop in new domains, numerous studies have concentrated on adapting DNNs trained on the source domain (w/ labels) to the target domain (w/o labels). Recent works in this field present significant advances even in challenging LiDAR-based 3D object detection. For example, [11, 12, 22] address the adaptation problem of Waymo [23] nuScenes [18], presenting notable improvement. Nevertheless, several drawbacks remain that (1) UDA often requires a sophisticated training strategy, e.g., adversarial training scheme [24, 22] and multiple rounds of self-training [25, 26]. (2) Potential risk of error propagation exists as many UDA methods rely on pseudo labeling [25, 26, 11, 12]. (3) Sensor data from the target domain, although not need to be labeled, are required which could be still expensive in practice.
II-C Synthetic Driving Datasets
Cost-efficient virtual environments have been actively employed to reduce the burden of constructing large-scale driving datasets with rich annotations. For example, Unity222http://unity3d.com/ engine is used to generate Virtual KITTI 2 (VKITTI2) [14], and SYNTHIA-AL [16]. A modern game Grand Theft Auto V is utilized to create GTA5 [19] and VIPER [13]. Recently, open-source driving simulator CARLA [27] has been employed to produce SHIFT [17].
Despite prevailing achievements of existing synthetic datasets, they are less suitable for scenarios where their role is to assist the adaptation of DNNs trained on to unobserved . The reason is that many synthetic datasets are generated from their own sensor suite and virtual worlds with no real-world references. Therefore, there is no guarantee that the geological features of , a target region for deploying AVs with DNNs, are present inside them. VKITTI2 is an exception as it is a digital-version clone of a real-world dataset, KITTI. The mentioned characteristics of VKITTI2 enable it to assess the usefulness of virtual worlds in terms of transferability to the real world. Nevertheless, this validation is limited to camera-related tasks, as LiDAR of KITTI is not implemented in VKITTI2. Furthermore, BBox labels for cyclists and pedestrians that exist in KITTI are absent, which makes it difficult to be used for learning about classes other than vehicles.
Our MORDA dataset benchmarks nuScenes, meaning it incorporates cameras, LiDAR, and 10 detection classes with accurate BBox labels for both modalities. Tab. I compares MORDA with publicly available synthetic datasets.
III MORDA DATASET GENERATION
III-A Overview
The purpose of our method is to construct a dataset from a synthetic-fusion domain () where DNNs (specifically, 2D/3D object detectors) can indirectly learn about the unforeseen to enhance adaptability while retaining the performance on at which the model is trained. To accomplish the aim, we extract pivotal characteristics from and . Then, they are reproduced and blended within a simulator, creating our fusion domain. Lastly, our synthetic dataset (MORDA) is generated using the simulator. Fig. 2 illustrates this procedure in detail. As a showcase of the proposed methodology, we choose nuScenes and South Korea to represent and , respectively. Next, the MORAI Simulator (MORAI SIM) [28] is employed to provide a baseline simulation environment where fundamental features like driving agents and GT-label generation are available.
III-B Reproduced Characteristics of Source Domain,
We extract three components from and implement them in MORAI SIM to create MORDA: sensor suite, list of object categories for detectors to classify, and labeling rules to draw 3D BBoxes for each category. As nuScenes is employed for , its sensor suite (six 1600900 cameras and one 32-channel spinning LiDAR), object categories (10 detection classes), and annotation rules for 3D BBox are reproduced. The positions and orientations of individual virtual sensors adhere to the configuration of real ones in nuScenes. It is worth mentioning that 3D BBox labeling rules even for complicated categories that have two individual rigid sections (e.g., bendy bus, truck-trailer) are implemented as visualized in Fig. 2. Further, the types of objects used for dataset generation include construction vehicles and road objects (barriers, traffic cones), which are rarely observed in other synthetic datasets at present.
III-C Reproduced Characteristics of Target Domain,
We want the geographical features of the real target domain, South Korea, to be reflected in the virtual environment so that our generated MORDA dataset could be enriched with them. To ensure this, we employ four digital-twin maps of South Korea available in MORAI SIM, which consist of one highway and three urban cities. Note that each digital-twin map was constructed upon HD map data, geographical survey data, and 3D building data collected from each counterparting region in South Korea so that the road surface, terrain, and surrounding static objects can be realistically replicated in the simulator.
III-D Data Creation in Synthetic Fusion Domain,
By reproducing the mentioned real-world characteristics in the simulator, we construct , the synthetic-fusion domain where we hypothesize that the features from and are appropriately blended. 87 scenes are made in this domain, and for each scene, the ego vehicle drives on a predefined path at a certain place of a digital twin map to collect synthetic sensor data. The scenes are composed of two types: static and dynamic, accounting for 57 and 30.
The major difference between static and dynamic scenes is the presence of moving objects. In static scenes, objects other than the ego vehicle are stationary. However, their arrangement has been meticulously handcrafted to emulate parking lots, construction sites, and road environments filled with huge vehicles. Hence, this type of scene has a larger number of trucks, trailers, construction vehicles, barriers, and traffic cones than the other type. In contrast to static scenes, the focus of dynamic scenes is to incorporate dense traffic flows with many moving objects into our dataset. To achieve this, we exploit the built-in traffic generator of MORAI SIM, which spawns various agents (e.g., cars and motorcyclists) driving autonomously, to keep the road congested.333The traffic flow was set to reach E (operation near or at capacity) or F (breakdown in flow) of Level-of-service (LOS). Details regarding this can be found in [29].
While having the ego vehicle drive in a scene, the synchronized virtual sensors collect data at 20 Hz along with ground-truth (GT) labels. Iterating this process across all scenes yields our MORDA dataset enriched with the simulated geographical features of and object-shape features of . Lastly and optionally, we convert MORDA to follow the dataset format of nuScenes. The lower-right corner of Fig. 2 illustrates our converted dataset visualized using nuScenes development toolkit (nuscenes-devkit) [18]. The reason for this conversion is detailed in Sec. V-A1
IV MORDA Dataset
IV-A Dataset Content
MORDA comprises ~37K frames where each frame consists of ego-vehicle pose data, six 1600900 images, and one 32-channel point cloud data with corresponding GT labels. The types of GT include pixel/point-level semantic and instance segmentation, 2D/3D BBox annotation, and pixel-wise depth values. All types of GT labels are generated by individual virtual sensors independently. Therefore, our dataset can be utilized for various vision tasks such as monocular/multi-view camera-based, lidar-based, fusion-based object detection, and segmentation tasks.
IV-B Comaprison to nuScenes Dataset
MORDA delivers 1.6 million (M) 3D BBox labels for ~37K frames, whereas nuScenes provides 40K frames with 1.4M BBoxes (only keyframes are counted excluding sweeps as they are not annotated). While both datasets present comparable scales of labeled frames, MORDA has ~15 more BBox labels than nuScenes. Furthermore, our dataset presents denser temporal information of objects around ego vehicles as MORDA’s BBox labels are generated at 20 Hz while those of nuScenes are labeled at 2 Hz.
Fig. 3 visualizes the count of 3D BBoxes for each class by dataset. As shown, their distributions largely differ, but MORDA supplements the shortage of 3D BBoxes for several classes in nuScenes. Specifically, the longtailness of rare classes (truck, bus, construction vehicle, trailer, motorcycle, and bicycles) in nuScenes is mitigated when 3D-BBox labels of respective classes in MORDA are concatenated.
Similarly, Fig. 4 depicts the distribution of 2D BBoxes in image space for each dataset where the individual 2D labels are represented as circular markers at coordinates corresponding to their respective widths and heights. In general, the two distributions show a similar tendency and coverage in the width-height 2D space. However, it is noteworthy that the distribution of our dataset shows higher density in the area highlighted by the red dotted ellipse where BBox labels with high values for height (500+) and a wide range for width (0 ~ 800) exist. This suggests that images in our dataset contain more diverse shapes of huge vehicles such as buses, construction vehicles, trucks, and trailers, likely at close distances. Combining all the above, mitigation of the class imbalance and enhancement of data diversity could be expected when adding MORDA to nuScenes.
[b] Task Modality Model Remark Training Dataset Evaluation Dataset nuScenes () AI-Hub () Truck Bus Trailer CV MC Bicycle mAP† NDS Car Truck Bus Ped mAP‡ NDS 2D C Faster-RCNN ✗ nuScenes 31.4 48.0 16.1 4.2 20.9 19.8 27.3 ✗ 19.5 13.8 16.8 3.3 13.35 ✗ 2D C Faster-RCNN ✗ nuScenes + MORDA 31.0 50.0 16.5 3.7 22.0 20.7 27.8 ✗ 27.2 18.6 24.5 8.3 19.7 ✗ –0.4 +2.0 +0.4 –0.5 +1.1 +0.9 +0.5 ✗ +7.7 +4.8 +7.7 +5.0 +6.35 ✗ 3D L PointPillars SECFPN nuScenes 39.2 51.7 28.6 5.4 21.5 0.9 34.94 50.02 24.93 3.425 11.70 10.93 12.74 21.95 3D L PointPillars SECFPN nuScenes + MORDA 40.5 53.1 30.8 7.3 22.9 0.8 35.45 50.44 29.67 5.59 12.28 11.99 14.88 22.63 +1.3 +1.4 +2.2 +1.9 +1.4 –0.1 +0.51 +0.42 +4.74 +2.165 +0.58 +1.06 +2.14 +0.68 3D L SSN ✗ nuScenes 49.5 65.8 33.7 17.1 52.6 23.3 48.29 59.42 27.30 9.93 19.94 15.48 18.16 26.92 3D L SSN ✗ nuScenes + MORDA 52.2 66.4 33.1 16.8 52.5 24.6 48.92 60.41 36.76 13.79 24.16 20.76 23.87 30.21 +2.7 +0.6 –0.6 –0.3 –0.1 +1.3 +0.63 +0.99 +9.46 +3.86 +4.22 +5.28 +5.71 +3.29 3D L CenterPoint Pillar (0.2m) nuScenes 49.0 63.3 31.4 10.9 41.4 18.6 48.96 59.42 0.0* 0.0* 0.065 0.0* 0.016 2.64 3D L CenterPoint Pillar (0.2m) nuScenes + MORDA 49.5 64.3 32.8 14.3 44.4 13.9 49.43 59.76 14.73 0.012 2.01 0.0* 4.19 11.38 +0.5 +1.0 +1.4 +3.4 +3.0 –4.7 +0.47 +0.34 +14.73 +0.012 +1.945 0.0 +4.17 +8.74 3D L CenterPoint Voxel (0.1m) nuScenes 53.2 66.5 36.0 15.0 55.2 36.8 56.23 64.51 0.717 0* 0* 0* 0.179 7.203 3D L CenterPoint Voxel (0.1m) nuScenes + MORDA 55.1 69.2 37.3 18.5 57.0 38.1 57.80 65.47 63.44 5.682 4.264 0* 18.34 22.93 +1.9 +2.7 +1.3 +3.5 +1.8 +1.3 +1.57 +0.96 +62.72 +5.682 +4.264 0.0 +18.16 +15.72 3D L CenterPoint Voxel (0.075m) nuScenes 55.1 67.9 34.9 15.2 55.8 36.1 56.95 65.40 0.197 0.0* 0.0* 0.0* 0.049 4.19 3D L CenterPoint Voxel (0.075m) nuScenes + MORDA 55.9 67.1 35.7 16.1 58.4 40.1 58.02 66.02 63.28 1.0 24.6 0.0* 22.22 23.71 +0.8 –0.8 +0.8 +0.9 +2.6 +4.0 +1.07 +0.62 +63.08 +1.0 +24.6 0.0 +22.17 +19.52
Over 10 detection classes of nuScenes.
Over 4 classes: car, truck, bus, and pedestrian.
If the precision or recall values for all operating points on the precision-recall curve are less than 10, the AP for that class is set to zero [18].
V EXPERIMENTS
V-A Impact of MORDA on (nuScenes dataset)
This section presents quantitative results that MORDA could maintain or even improve the performance of camera-based 2D and LiDAR-based 3D detectors on nuScenes.
V-A1 Pre-processing
We convert MORDA dataset to follow the format of nuScenes (e.g., structure, file hierarchy, and keyframe-sweep), primarily meaning annotation frequency of MORDA decrease from 20 Hz to 2 Hz. This step is necessary as numerous modern 3D detectors trained on nuScenes, especially LiDAR-based ones, exploit the sweeps (frames w/o annotations) from past time steps for improving point cloud density and robustness of temporal information. Thus, to construct a coherent training environment for those detectors we conduct the format conversion although it decreases the frequency of MORDA (keyframes: ~3.7K, sweeps: ~33K).
V-A2 Training details
No advanced training or domain-adaptation strategy is applied other than a simple concatenation of nuScenes and MORDA to construct a merged training set. For monocular 2D detection scenario, Faster-RCNN [30] is opted. We employ the implementation of Faster-RCNN from MMDetection [31], and the default training configuration (e.g., learning rate, batch size) remains unchanged. The detector is trained on one RTX 3090 GPU with images that are from the front camera. For LiDAR-based 3D detection, renowned PointPillars [4], SSN [5], and CenterPoint [6] are employed. we use the implementation of the mentioned networks in MMDetection3D [32]. While the other configurations are the same as the default ones, batch size is adjusted to our GPU resources. In short, PointPillars, and SSN are trained on a single RTX 3090 with batch size of 8. CenterPoint is trained on four NVIDIA L4 GPUs with batch sizes set to 8, 8, and 4 for Pillar (0.2m) and Voxel (0.1m, 0.075m), respectively.
V-A3 Metrics
We use renowned average precision (AP), mean average precision (mAP), and nuScenes detection score (NDS) [18]. All metrics are represented in percentage (), and the higher the better.
V-A4 Evaluation
The left half of Tab. II presents performance of detectors on the validation split of nuScenes. As demonstrated, detectors trained on nuScenes combined with MORDA outperform all baselines that are trained on nuScenes alone in terms of mAP and NDS across modalities and network architectures. While the mAP gains achieved in the other detectors are marginal, MORDA brings notable improvement of 1+ to CenterPoint-Voxel (0.1m, 0.075m) models. This gain is worth mentioning as no tailored training strategy was applied to close the sim-to-real gap explicitly.
In addition to mAP, we present class-wise AP for the rare categories that MORDA mitigates long-tail issues discussed in Sec. IV-B. Although the trend of AP improvement or decline varies across models for all six classes, we observe meaningful enhancements, e.g., +2.7 for truck in SSN, +3.4 and +3.0 for CV and MC in CenterPoint-Pillar, and +4.0 for bicycle in CenterPoint-Voxel(0.075m). Combining all the above leads us to conclude the characteristics of nuScenes we benchmark are well implemented in MORDA, and it does not degrade detectors’ performance on nuScenes overall.
V-B Impact of MORDA on (AI-Hub dataset)
To validate the efficacy of simulated features provided by MORDA in terms of previewing and indirectly learning about , we load the trained models from Sec. V-A and evaluate their performance on without further training.
T and V denote Train and Validation, respectively.
[b] Annot. Frame CAM. Resolution Beam Det. Class Freq. (Hz) nuScenes (T) 28K 61 1600900 32 10 2 MORDA (T) 37K3.7K 6 1 1600900 32 10 20 2 AI-Hub (V) 80K8.8K 5 19201200 1600900 128 84 102
V-B1 Real-target dataset
A real-world driving dataset, AI-Hub is opted to represent given (1) its sensor data are collected from the real-target domain, South Korea where digital twin maps in MORDA () is originated from. (2) Sensor suite and GT types are similar to nuScenes and MORDA, i.e., each frame contains five 19201200 multi-view images, one point cloud from 128-channel spinning LiDAR, and RTK GNSS data with 3D-BBox annotation frequency of 10 Hz.
V-B2 Pre-processing
Tab. III summarizes the applied modifications. Format of AI-Hub dataset is converted to nuScenes style, following Sec. V-A1. On top of that, we apply additional processing for fair experiments with Sec. V-A. (1) Only images from the front camera are used for 2D detection, and they are cropped to 1600900. (2) We generated pseudo-2D-BBox labels exploiting pixel-level panoptic segmentation GT as they are not present in AI-Hub. Lastly, (3) we take 3D BBoxes only for four classes out of eight, i.e., car, truck, bus, and pedestrian, as they either have deviating labeling policies from nuScenes or have few instances ( 100). Note that all data of AI-Hub are used for validation, not training.
V-B3 Evaluation
The right half of Tab. II shows detection scores of models evaluated on the unobserved AI-Hub dataset. As can be seen, detectors trained with MORDA exceed respective baselines by a large margin across modalities. For example, mAP gains of 6.3 for camera-based Faster-RCNN and 5.7 for LiDAR-based SSN are reported. This suggests that MORDA can provide simulated real-world experiences, assisting the adaptation of detectors and lessening reliance on real sensor data from .
In addition to the mentioned benefit above, we observe MORDA can assist in mitigating the generalization failure. Tab. II shows that the baseline CenterPoint family reports severe performance degradation under domain shift (nuScenesAI-Hub), meaning their class-wise AP and mAP scores on AI-Hub almost reach to zero regardless of backbone types. In contrast, when MORDA is included, mAP scores of all CenterPoint models get improved, meaning the observed generalization failures get alleviated. Especially, those with voxel backbones (0.1m and 0.075m) gain significant improvements in mAP (+18 and +22). Fig. 5 explores this further and illustrates the development of mAP throughout entire training process with respect to the presence of MORDA. Notice that the model with MORDA achieves a substantial performance gap compared to the baseline on AI-Hub. In light of the analysis above, we conclude that MORDA is a well-constructed preview for and a useful regularization tool for generalization failure. Fig. 6 illustrates benefits of MORDA in this aspect.
[b] Training Dataset Frame mAP (, ) nuScenes 28K 27.3 nuScenes + VKITTI2 28K + 2.5K 27.0 (–0.3) nuScenes + SYNTHIA-AL 28K + 7.3K 27.4 (+0.1) nuScenes + SHIFT 28K + 30K 27.7 (+0.4) nuScenes + MORDA (Ours) 28K + 3.7K 27.8 (+0.5)
V-C Benchmark with Other Synthetic Datasets: 2D Detection
VKITTI2, SYNTHIA-AL, and SHIFT datasets are compared to MORDA in terms of mAP gains that respective datasets bring to Faster-RCNN trained on nuScenes.
V-C1 Pre-processing
We use images from the front camera that all mentioned datasets above have in common. Next, we sample the datasets to match their frequencies as much as possible. That means MORDA, SYNTHIA-AL, and VKITTI2444The frequency of VKITTI2 is assumed to be the same as KITTI, as we could not find relevant information. are downsampled to 2Hz. Although images of SHIFT dataset (discrete-shift) are already of 1 Hz, we downsample it by a factor of 5 due to its exceptional size (150K). Note that SHIFT is still the largest dataset after this downsampling, as shown in Tab. IV. 2D BBoxes of respective datasets are merged to nuScenes by class labels.
V-C2 Training Details
Faster-RCNN is trained on nuScenes and either of the synthetic datasets for all 10 detection classes, using MMDetection with default configuration.
V-C3 Evaluation
Tab. IV presents mAPs of Faster-RCNNs trained with respective synthetic datasets along with the numbers of frames used for training. Our MORDA achieves 27.8, surpassing the baseline and all the compared datasets. SHIFT shows a comparable mAP gain with ours, but it used 8 more frames than MORDA.
VI CONCLUSIONS
For safety-critical AVs, DNNs need to overcome the performance degradation that occurs in unexperienced while maintaining performance on . This paper proposed leveraging where key characteristics of and are fused. To showcase the efficacy, we generated a novel dataset, MORDA. Our comprehensive experiments demonstrate that MORDA could help DNNs adapt to , mitigating their dependency on real-world data without losing performance on . Nevertheless, the potential of MORDA is not fully explored as only the training strategy of simple dataset concatenation was applied. Thus, future work could be investigating sophisticated training methods to further enhance the performance on .